OpenStudy (mrnebula):

HELP MEDAL!!! What is the solution of the system? 7x + 5y = 19 -7x - 2y = -16 I know that I am suppose to use elimination to solve it but I dont understand the example in my text book. HELP!

undeadknight26 (undeadknight26):

omg im doing this today >.>

OpenStudy (mrnebula):

lol

OpenStudy (mrnebula):

there is an example in my textbook but no help

OpenStudy (aripotta):

yes, the elimination method would be the best way to go since it's already set up to eliminate x. first, combine the two equations

OpenStudy (mrnebula):

how?

syed98 (syedmohammed98):

7x + 5y=19 x = (19-5y)/7 -7((19-5y)/7) + 5y= 19 solve it ..u get the value of y simply insert the value of y in 7x + 5y=19 to fine the value of x :)

OpenStudy (aripotta):